### v3

# Convergence rates for polynomial optimization on set products

Abstract
We consider polynomial optimization problems on Cartesian products of basic compact semialgebraic sets. The solution of such problems can be approximated as closely as desired by hierarchies of semidefinite programming relaxations, based on classical sums of squares certificates due to Putinar and Schmüdgen. When the feasible set is the bi-sphere, i.e., the Cartesian product of two unit spheres, we show that the hierarchies based on the Schmüdgen-type certificates converge to the global minimum of the objective polynomial at a rate in $O(1/t^{2})$, where $t$ is the relaxation order. Our proof is based on the polynomial kernel method. We extend this result to arbitrary sphere products and give a general recipe to obtain convergence rates for polynomial optimization over products of distinct sets. Eventually, we rely on our results for the bi-sphere to analyze the speed of convergence of a semidefinite programming hierarchy approximating the order $2$ quantum Wasserstein distance.
Keywords. polynomial optimization, Cartesian product, Moment-SOS hierarchy, Positivstellensatz, polynomial kernel method, quantum Wassertein distance
1 Introduction
Given a tuple $\mathbf{x}=(x_{1},\dots,x_{n})$, we denote by $\mathbb{R}[\mathbf{x}]$ the vector space of polynomials with real coefficients. Given $r\in\mathbb{N}$ and a finite set of polynomials $q,g_{1},\dots,g_{r}\subset\mathbb{R}[\mathbf{x}]$, we consider the problem of minimizing $q$ over the basic compact semialgebraic set
$\mathbf{X}:=\{\mathbf{x}\in\mathbb{R}[\mathbf{x}]:g_{1}(\mathbf{x})\geq 0,\dots,g_{r}(\mathbf{x})\geq 0\},$
namely to compute $q_{\min}:=\inf_{\mathbf{x}\in\mathbf{X}}q(\mathbf{x})$. Note that this polynomial optimization problem is equivalent to
$q_{\min}=\sup_{\tau\in\mathbb{R}}\{\tau:q-\tau\in\mathcal{P}_{+}(\mathbf{X})\},$ (1)
where $\mathcal{P}_{+}(\mathbf{X})$ is the set of polynomials being nonnegative on $\mathbf{X}$. Polynomial optimization problems of the form (1) capture difficult instances such as the NP-hard MaxCut problem. We refer the interested reader to the recent monographs [HKL20, Nie23, MW23] that address several theoretical and applicative aspects of polynomial optimization.
Since the characterization of $\mathcal{P}_{+}(\mathbf{X})$ is a highly non-trivial task, there have been substantial research efforts to provide tractable inner approximations, based on sums of squares decompositions. A polynomial $p\in\mathbb{R}[\mathbf{x}]$ is called a sum of squares if it can be decomposed as $p=\sum_{i}q_{i}^{2}$, for some $q_{i}\in\mathbb{R}[\mathbf{x}]$. The set of sums of squares of polynomials is denoted by $\Sigma[\mathbf{x}]$.
Let us define $g_{0}:=1$, $[r]:=\{1,\dots,r\}$ and $g_{J}:=\displaystyle\prod_{j\in J}g_{j}$, for all $J\subseteq[r]$.
For instance one can approximate $\mathcal{P}_{+}(\mathbf{X})$ from the inside while relying either on the preordering $\mathcal{T}(\mathbf{X})\subseteq\mathcal{P}_{+}(\mathbf{X})$ or the quadratic module $\mathcal{Q}(\mathbf{X})\subseteq\mathcal{P}_{+}(\mathbf{X})$, defined as
$\mathcal{T}(\mathbf{X})$ :=$:=\left\{\sum_{J\in[r]}\sigma_{J}g_{J}:\sigma_{J}\in\Sigma[\mathbf{x}]\right\},$ (2)
$\mathcal{Q}(\mathbf{X})$ :=$:=\left\{\sum_{j=0}^{m}\sigma_{j}g_{j}:\sigma_{j}\in\Sigma[\mathbf{x}]\right\}.$ (3)
When $\mathbf{X}$ is compact, Schmüdgen’s Positivstellensatz states that every polynomial positive on $\mathbf{X}$ belongs to $\mathcal{T}(\mathbf{X})$:
Theorem 1 (Schmüdgen’s Positivstellensatz [Sch91]).
Let $\mathbf{X}$ be a basic compact semialgebraic set. Then for any polynomial $q\in\mathcal{P}_{+}(\mathbf{X})$ and $\varepsilon>0$, one has $q+\varepsilon\in\mathcal{T}(\mathbf{X})$.
Putinar’s Positivstellensatz allows one to replace $\mathcal{T}(\mathbf{X})$ by $\mathcal{Q}(\mathbf{X})$ in the above result under an assumption slightly stronger than compactness. Given a vector $\mathbf{x}\in\mathbb{R}^{n}$, its Euclidean norm is denoted by $\|\mathbf{x}\|:=\sqrt{\sum_{i=1}^{n}x_{i}^{2}}$. The quadratic module $\mathcal{Q}(\mathbf{X})$ is called Archimedean if $N-\|\mathbf{x}\|^{2}\in\mathcal{Q}(\mathbf{X})$.
Theorem 2 (Putinar’s Positivstellensatz [Put93]).
Let $\mathbf{X}$ be a basic semialgebraic set and assume that $\mathcal{Q}(\mathbf{X})$ is Archimedean. Then for any polynomial $q\in\mathcal{P}_{+}(\mathbf{X})$ and $\varepsilon>0$, one has $q+\varepsilon\in\mathcal{Q}(\mathbf{X})$.
1.1 Hierarchies of lower bounds for polynomial optimization
For a given integer $t\geq\max_{1\leq j\leq r}\{\lceil\deg g_{j}/2\rceil\}$, if one considers the truncated preordering or quadratic module defined as
$\mathcal{T}(\mathbf{X})_{2t}$ :=$:=\left\{\sum_{J\in[r]}\sigma_{J}g_{J}:\sigma_{J}\in\Sigma[\mathbf{x}],\quad\deg(\sigma_{J}g_{J})\leq 2t\right\},$ (4)
$\mathcal{Q}(\mathbf{X})_{2t}$ :=$:=\left\{\sum_{j=0}^{m}\sigma_{j}g_{j}:\sigma_{j}\in\Sigma[\mathbf{x}],\quad\deg(\sigma_{j}g_{j})\leq 2t\right\},$ (5)
then checking that a polynomial belongs to either $\mathcal{T}(\mathbf{X})_{2t}$ or $\mathcal{Q}(\mathbf{X})_{2t}$ can be done by solving a semidefinite program. Semidefinite programming is a particular class of convex programming, that consists of minimizing a linear function under linear matrix inequality constraints [VB96], and for which dedicated solvers can be readily used.
It follows from the work by Lasserre [Las01] that the hierarchies of Schmüdgen-type and Putinar-type lower bounds defined as
$\operatorname{lb}(q,\mathcal{T}(\mathbf{X}))_{t}:=\sup_{\tau\in\mathbb{R}}\{\tau:q-\tau\in\mathcal{T}(\mathbf{X})_{2t}\},$ (6)
$\operatorname{lb}(q,\mathcal{Q}(\mathbf{X}))_{t}:=\sup_{\tau\in\mathbb{R}}\{\tau:q-\tau\in\mathcal{Q}(\mathbf{X})_{2t}\}.$ (7)
both converge to $q_{\min}$ when $t$ goes to infinity, as a consequence of Schmüdgen’s and Putinar’s Positivstellensatz, respectively. Note that when $\mathbf{X}$ involves either equality constraints only or a single inequality constraint, then the preordering and quadratic module coincide. However, only the former remains true when one considers Cartesian products.
1.2 Related works on convergence rates
Recently, there has been a lot of dedicated research efforts in analyzing the asymptotic behavior of the hierarchies of Schmüdgen-type and Putinar-type bounds defined in (6)-(7), respectively. For a polynomial $q$ achieving its minimal value $q_{\min}$ over a general compact semialgebraic set $\mathbf{X}$, it was proved in [Sch04] that the Schmüdgen-type bounds $\operatorname{lb}(q,\mathcal{T}(\mathbf{X}))_{t}$ converge to $q_{\min}$ at a rate in $O(1/t^{c})$, where $c$ is a constant depending on $\mathbf{X}$. If $\mathcal{Q}(\mathbf{X})$ is Archimedean, an initial result provided in [NS07] shows that the Putinar-type bounds $\operatorname{lb}(q,\mathcal{Q}(\mathbf{X}))_{t}$ converge to $q_{\min}$ at a rate in $O(1/\log(t)^{c})$. Recently, this rate has been exponentially improved in [BM23], where the authors prove that the Putinar-type bounds converge to $q_{\min}$ at a rate in $O(1/t^{c})$, thus matching the convergence rate of Schmüdgen-type bounds. Without any compactness assumption on the feasible set, one can rely on Putinar-type hierarchies based on sums of squares of rational fractions with (prescribed) uniform denominators. For these hierarchies, a similar rate in $O(1/t^{c})$ has been obtained in [MM22].
For special cases of feasible sets including the unit sphere/ball, the standard simplex and the hypercube, the value of the constant $c$ can be explicitly given. An improved convergence rate in $O(1/t^{2})$ for Schmüdgen-type bounds has been obtained in [FF21] for the unit sphere, in [Slo22] for the unit ball and the standard simplex, and in [LS23] for the hypercube. The case of the binary hypercube $\{0,1\}^{n}$ is covered in [SL23]. For Putinar-type bounds, a rate in $O(1/t)$ has been recently obtained in [BS24] for the hypercube. For polynomial optimization with correlative sparsity, stronger rates have been obtained in [KMRZ25] under the assumption that the input data are (sufficiently) sparse.
In addition, convergence rates are available for the generalized moment problem (GMP), which consists of minimizing a linear function over the cone of positive Borel measures supported on $\mathbb{R}^{n}$. Note that polynomial optimization can be cast as a special instance of the GMP, and that Schmüdgen-type and Putinar-type hierarchies can be naturally derived for other GMP instances. When the GMP instance involves finitely many constraints, one can obtain a rate similar to the one for polynomial optimization on the measure support [GM25]. Specialized rates have been obtained for certain GMP instances involving infinitely (countably) many constraints, in particular for optimal control [KHJ17], volume estimation [KH18], and exit location estimation of stochastic processes [ST24].
We refer to the recent survey [LS24] for a more detailed overview of convergence rates and related aspects, in particular the tightness of the performance analysis, the exponential convergence under local optimality conditions, and finite convergence.
1.3 Our contributions
In this work, we mainly focus on hierarchies of Schmüdgen-type bounds for polynomial optimization over Cartesian products of compact semialgebraic sets.
We first provide in Section 3.1 a convergence rate of Schmüdgen-type (or equivalently Putinar-type) bounds when minimizing a polynomial on the product of spheres $\mathbf{X}=S^{n-1}\times S^{n-1}$. Given two $n$-dimensional tuples of real variables $\mathbf{x}=(x_{1},\dots,x_{n})$ and $\mathbf{y}=(y_{1},\dots,y_{n})$, and a polynomial $q\in\mathbb{R}[\mathbf{x},\mathbf{y}]$, we define $q_{\min}:=\min_{(\mathbf{x},\mathbf{y})\in\mathbf{X}}q(\mathbf{x},\mathbf{y})=\min_{\mathbf{x},\mathbf{y}\in S^{n-1}}q(\mathbf{x},\mathbf{y})$ and $q_{\max}:=\max_{(\mathbf{x},\mathbf{y})\in\mathbf{X}}q(\mathbf{x},\mathbf{y})$.
Theorem 3.
Let $\mathbf{X}=S^{n-1}\times S^{n-1}$ be the Cartesian product of two unit spheres, and let $q\in\mathbb{R}[\mathbf{x},\mathbf{y}]$ be a polynomial of degree $d$. Then for any $t\geq 2nd\sqrt{d}$, the lower bound $\operatorname{lb}(q,\mathcal{T}(\mathbf{X}))_{2t}$ for the minimization of $q$ over $\mathbf{X}$ satisfies:
$q_{\min}-\operatorname{lb}(q,\mathcal{T}(\mathbf{X}))_{2t}\leq\frac{C_{\mathbf{X}}(n,d)}{t^{2}}\cdot(q_{\max}-q_{\min}).$ (8)
Here, $C_{\mathbf{X}}(n,d)$ is a constant depending only on $n$ and $d$. This constant has a polynomial dependence on $n$ at fixed $d$, and a polynomial dependence on $d$ at fixed $n$. See relation (23) for details.
The proof of Theorem 3 is given in Section 3.1.
We extend this result in Section 3.2 (Theorem 8) to the case of arbitrary sphere products, and in Section 3.3 (Theorem 9) to the case of products of distinct sets.
Eventually we analyze in Section 4 the convergence rate of a problem arising from quantum information theory, called the quantum Wasserstein distance. Computing this distance boils down to solving a specific instance of the generalized moment problem on the complex bi-sphere. We provide a rate in $O(1/t^{2})$, stated in Theorem 11, by first changing variables from complex to real, then by combining Theorem 3 with the recent convergence rate results from [GM25], obtained for the generalized moment problem.
1.4 Outline of the proof technique
For a given compact semialgebraic set $\mathbf{X}\subset\mathbb{R}^{n}$ and a polynomial $p\in\mathbb{R}[\mathbf{x}]$, we denote the supremum-norm of $p$ on $\mathbf{X}$ by $\|p\|_{\mathbf{X}}:=\max_{\mathbf{x}\in\mathbf{X}}|p(\mathbf{x})|$. To obtain the convergence rate of Theorem 3, we follow the exact same technical strategy pursued in [FF21, Slo22] for the case of the unit sphere $S^{n-1}$, the unit ball and the standard simplex, as well as in [LS23] for the case of the hypercube. Let us denote by $\mathbb{R}[\mathbf{x},\mathbf{y}]_{d}$ the vector space of polynomials of degree at most $d$. Similarly let $\mathcal{P}_{+}(\mathbf{X})_{d}$ be the subset of polynomials from $\mathbb{R}[\mathbf{x},\mathbf{y}]_{d}$ that are nonnegative on $\mathbf{X}$. Let $q\in\mathbb{R}[\mathbf{x},\mathbf{y}]_{d}$ and $\mathbf{X}=S^{n-1}\times S^{n-1}$. The goal is to prove that
$q-q_{\min}+\varepsilon\in\mathcal{T}(\mathbf{X})_{4t}$
for some small $\varepsilon$. Up to translation and scaling, we assume that $q_{\min}=0$ and $q_{\max}=1$. We will construct an invertible linear operator $\operatorname{\mathbf{K}}:\mathbb{R}[\mathbf{x},\mathbf{y}]_{d}\to\mathbb{R}[\mathbf{x},\mathbf{y}]_{d}$ which satisfies the following properties:
$\operatorname{\mathbf{K}}(1)=1,$ (P1)
$\operatorname{\mathbf{K}}p\in\mathcal{T}(\mathbf{X})_{4t},\quad\text{for all }p\in\mathcal{P}_{+}(\mathbf{X})_{d},$ (P2)
$\|\operatorname{\mathbf{K}}^{-1}q-q\|_{\mathbf{X}}\leq\varepsilon.$ (P3)
Then it follows from [Slo22, Lemma 7] that $q+\varepsilon\in\mathcal{T}(\mathbf{X})_{4t}$ and thus $q_{\min}-\operatorname{lb}(q,\mathcal{T}(\mathbf{X}))_{2t}\leq\varepsilon$. As in [FF21, Slo22, LS23], the statement of Theorem 3 is proven by showing the existence, for each large enough $t$, of an operator $\operatorname{\mathbf{K}}$ satisfying (P1), (P2) and (P3) with $\varepsilon=O(1/t^{2})$.
The way to construct such an operator is to follow the so-called polynomial kernel method. In Section 2, we explain how such a kernel is found in [FF21] for the case of the unit sphere. Our rate for the bi-sphere is obtained by multiplying the two kernels associated to each unit sphere. This idea of kernel multiplication is the same as the one previously used for the hypercube in [LS23] (see also [KMRZ25]).
2 Preliminaries: convergence rate on the unit sphere
For a multi-index $\alpha=(\alpha_{1},\dots,\alpha_{n})\in\mathbb{N}^{n}$, we use the notation $|\alpha|=\alpha_{1}+\dots+\alpha_{n}$. Given $d\in\mathbb{N}$, let $\mathbb{N}_{d}^{n}:=\{\alpha\in\mathbb{N}^{n}:|\alpha|\leq d\}$. The dimension of $\mathbb{N}_{d}^{n}$ is equal to the number of monomials of degree at most $d$, i.e., $\binom{n+d}{d}$.
A map $\operatorname{C}:S^{n-1}\times S^{n-1}:\mathbb{R}$ on $S^{n-1}$ is called a polynomial kernel when $\operatorname{C}(\mathbf{x},\mathbf{x}^{\prime})$ is a polynomial in the variables $\mathbf{x}$, $\mathbf{x}^{\prime}$. Let us equip $S^{n-1}$ with the uniform Haar measure $\mu$, and consider the Christoffel-Darboux kernel $\operatorname{C}_{2t}$, which is defined in terms of an orthonormal basis $\{P_{\alpha}:\alpha\in\mathbb{N}^{n}\}$ for $\mathbb{R}[\mathbf{x}]$ with respect to $(S^{n-1},\mu)$ as:
$\operatorname{C}_{2t}(\mathbf{x},\mathbf{x}^{\prime}):=\sum_{k=0}^{2t}\operatorname{C}^{(k)}(\mathbf{x},\mathbf{x}^{\prime}),\quad\text{where }\operatorname{C}^{(k)}(\mathbf{x},\mathbf{x}^{\prime}):=\sum_{|\alpha|=k}P_{\alpha}(\mathbf{x})P_{\alpha}(\mathbf{x}^{\prime}).$ (9)
Then one associates to $\operatorname{C}_{2t}$ an operator $\operatorname{\mathbf{C}}_{2t}:\mathbb{R}[\mathbf{x}]\to\mathbb{R}[\mathbf{x}]$, defined as follows:
$\operatorname{\mathbf{C}}_{2t}p(\mathbf{x}):=\int_{S^{n-1}}\operatorname{C}_{2t}(\mathbf{x},\mathbf{x}^{\prime})p(\mathbf{x}^{\prime})\operatorname{d}\mu(\mathbf{x}^{\prime}).$ (10)
This operator $\operatorname{\mathbf{C}}_{2t}$ is called reproducing as it satisfies $\operatorname{\mathbf{C}}_{2t}p=p$, for all $p\in\mathbb{R}[\mathbf{x}]_{2t}$. Therefore, all its eigenvalues are equal to $1$.
For each $k$, the set $H_{k}[\mathbf{x}]:=\operatorname{span}\{P_{\alpha}:\alpha\in\mathbb{N}^{n},|\alpha|=k\}$ is the spanning set of orthonormal spherical harmonics of degree $k$, i.e., homogeneous polynomials of degree $k$ lying in the kernel of the Laplace operator. One has $p(\mathbf{x})=\sum_{k=0}^{d}p_{k}(\mathbf{x})$, for all $\mathbf{x}\in S^{n-1}$, with $p_{k}\in H_{k}$ given by
$p_{k}=\int_{S^{n-1}}\operatorname{C}^{(k)}(\mathbf{x},\mathbf{x}^{\prime})p(\mathbf{x}^{\prime})\operatorname{d}\mu(\mathbf{x}^{\prime}).$
In this case one has the following closed form of the Christoffel-Darboux kernel
$\operatorname{C}_{2t}(\mathbf{x},\mathbf{x}^{\prime})=\sum_{k=0}^{2t}\mathcal{G}_{k}^{(n-1)}(\mathbf{x}\cdot\mathbf{x}^{\prime}),\quad\text{forall }\mathbf{x},\mathbf{x}^{\prime}\in S^{n-1}.$ (11)
In the above formula, the set $\{\mathcal{G}_{k}^{(n)}:k\in\mathbb{N}\}$ of Gegenbauer polynomials are defined, for all $n\geq 2$, as the set of orthogonal polynomials with respect to the weight function $w_{n}(x):=\hat{w}_{n}(1-x^{2})^{\frac{n-2}{2}}$, where $\hat{w}_{n}$ is a normalization constant ensuring that $\int_{-1}^{1}w_{n}(x)\operatorname{d}x=1$. We refer the interested reader to [Sze75] for a detailed survey about univariate orthogonal polynomials. The formula (11) is also known as the Funk-Hecke formula, stated for instance in [Rub15, Theorem A.34].
The next step is to perturb the eigenvalues of the reproducing operator $\operatorname{\mathbf{C}}_{2t}$ by considering the following kernel instead:
$\operatorname{C}_{2t}(\mathbf{x},\mathbf{x}^{\prime};\lambda):=\sum_{k=0}^{2t}\lambda_{k}\operatorname{C}^{(k)}(\mathbf{x},\mathbf{x}^{\prime})=\sum_{k=0}^{2t}\lambda_{k}\mathcal{G}_{k}^{(n-1)}(\mathbf{x}\cdot\mathbf{x}^{\prime}),\quad\text{for all }\mathbf{x},\mathbf{x}^{\prime}\in S^{n-1}.$ (12)
Then $\lambda=(\lambda_{k})$ is the set of eigenvalues of the operator obtained in (10) by considering $\operatorname{C}_{2t}(\mathbf{x},\mathbf{x}^{\prime};\lambda)$ instead of $\operatorname{C}_{2t}(\mathbf{x},\mathbf{x}^{\prime})$. The following result provides one way to prove that $\operatorname{\mathbf{C}}_{2t}$ satisfies (P2).
Lemma 4 (See [Slo22, Lemma 11]).
Let $\mathbf{X}$ be a compact semialgebraic set, and let $\mu$ be a finite measure supported on $\mathbf{X}$. Let $Q$ be a convex cone, and suppose that $\operatorname{C}:\mathbf{X}\times\mathbf{X}\to\mathbb{R}$ is a polynomial kernel for which $\operatorname{C}(\cdot,\mathbf{x}^{\prime})\in Q$ for each $\mathbf{x}^{\prime}\in\mathbf{X}$ fixed. Then if $p\in\mathbb{R}[\mathbf{x}]$ is nonnegative on $\mathbf{X}$, we have $\operatorname{\mathbf{C}}p\in Q$. That is, when selecting $Q=\mathcal{T}(\mathbf{X})_{2t}$, the operator $\operatorname{\mathbf{C}}$ satisfies (P2).
The proof of Lemma 4 relies on Tchakaloff’s Theorem [Tch57] and the exisence of cubature rules. Now assume that $\lambda=(\lambda_{k})$ is the set of coefficients of a univariate sum of squares in the basis of Gegenbauer polynomials. Then at each fixed $\mathbf{x}^{\prime}$, the polynomial $\operatorname{C}_{2t}(\mathbf{x},\mathbf{x}^{\prime};\lambda)$, defined in (12), is equal to a sum of squares of degree at most $2t$ on $\mathbf{X}$, so $\operatorname{C}_{2t}(\mathbf{x},\mathbf{x}^{\prime};\lambda)\in\mathcal{T}(\mathbf{X})_{2t}$ and Lemma 4 implies that $\operatorname{\mathbf{C}}_{2t}$ satisfies (P2).
Next, if the harmonic decomposition of $q$ is $q(\mathbf{x})=\sum_{k=0}^{d}q_{k}(\mathbf{x})$ for all $\mathbf{x}\in S^{n-1}$, with $q_{k}\in H_{k}[\mathbf{x}]$, then one has $\operatorname{\mathbf{C}}_{2t}q=\sum_{k=0}^{d}\lambda_{k}q_{k}$. If in addition, $\lambda_{0}=1$ and $\lambda_{k}\neq 0$ for all $k$ then $\operatorname{\mathbf{C}}_{2t}^{-1}q=\sum_{k=0}^{d}1/\lambda_{k}q_{k}$. By [Slo22, Lemma 14], the resulting perturbed operator satisfies (P1) if $\lambda_{0}=1$, and it satisfies (P3) with
$\varepsilon=\max_{1\leq k\leq d}\max_{\|\mathbf{x}\|=1}|q_{k}(\mathbf{x})|\cdot\left|1-1/\lambda_{k}\right|.$
It remains to give an upper bound of the two multiplied quantities. For the first quantity, one can define the so-called harmonic constant:
$\gamma(S^{n-1})_{d}:=\max_{p\in\mathbb{R}[\mathbf{x}]_{d}}\max_{0\leq k\leq d}\frac{\|p_{k}\|_{S^{n-1}}}{\|p\|_{S^{n-1}}}.$ (13)
In case of minimizing a homogeneous polynomial of degree $d$, one can define a similar constant denoted by $\gamma(S^{n-1})_{=d}$. This latter constant can be bounded independently of the dimension $n$ as a consequence of [FF21, Proposition 5]. More generally, one has the following result.
Lemma 5.
For all integers $d,n\geq 3$, the constant $\gamma(S^{n-1})_{d}$ depends polynomially on $n$ (when $d$ is fixed) and polynomially on $d$ (when $n$ is fixed). One has
$\gamma(S^{n-1})^{2}_{d}\leq\max_{0\leq k\leq d}\mathcal{G}_{k}^{n}(1)=\max_{0\leq k\leq d}\left(1+\frac{2k}{n-2}\right)\cdot\binom{k+n-3}{k}.$ (14)
Proof.
As in the proof of [Slo22, Proposition 15], one relies on the reproducing properties of the Christoffel-Darboux kernel to show that
$\gamma(S^{n-1})^{2}_{d}\leq\max_{0\leq k\leq d}\max_{\mathbf{x}\in S^{n-1}}\operatorname{C}^{(k)}(\mathbf{x},\mathbf{x})=\max_{0\leq k\leq d}\mathcal{G}_{k}^{(n-1)}(1).$
The desired expression then follows from the value of the Gegenbauer polynomials at $1$, see, e.g., (2.9) in [Xu98]. ∎
To finally obtain the desired convergence rate, the second quantity can be bounded as follows.
Lemma 6 (See [FF21, Theorem 6]).
Let $n,d\in\mathbb{N}$. Then for every $t\geq 2nd\sqrt{d}$ there exists a univariate sum of squares $\sigma(x)=\sum_{k=0}^{2t}\lambda_{k}\mathcal{G}_{k}^{(n)}(x)$ of degree $2t$ with $\lambda_{0}=1$, $1/2\leq\lambda_{k}\leq 1$, for all $1\leq k\leq d$, and
$\sum_{k=1}^{d}(1-\lambda_{k})$ $\leq\frac{n^{2}d^{3}}{t^{2}},$ (15)
$\sum_{k=1}^{d}\left|1-1/\lambda_{k}\right|\leq 2\sum_{k=1}^{d}(1-\lambda_{k})$ $\leq\frac{2n^{2}d^{3}}{t^{2}}.$
3 Convergence rate for polynomial optimization on set products
3.1 Products of two spheres
This section is dedicated to the proof of Theorem 3.
Proof.
Up to translation and scaling, we assume that $q_{\min}=0$ and $q_{\max}=1$. Let us equip $\mathbf{X}=S^{n-1}\times S^{n-1}$ with the product measure $\mu_{\mathbf{X}}:=\mu\otimes\mu$, where $\mu$ is the uniform Haar measure on $S^{n-1}$. Given the orthonormal basis $\{P_{\alpha}:\alpha\in\mathbb{N}^{n}\}$ for $\mathbb{R}[\mathbf{x}]$ (respectively for $\mathbb{R}[\mathbf{y}]$) with respect to $(S^{n-1},\mu)$, $\{P_{\alpha}(\mathbf{x})P_{\beta}(\mathbf{y}):\alpha,\beta\in\mathbb{N}^{n}\}$ is an orthonormal basis for $\mathbb{R}[\mathbf{x},\mathbf{y}]$ with respect to $(\mathbf{X},\mu_{\mathbf{X}})$. With $t\geq 2nd\sqrt{d}$, and $\lambda=(\lambda_{k})$ as in Lemma 6, we build the desired operator $\operatorname{\mathbf{K}}_{2t}:\mathbb{R}[\mathbf{x},\mathbf{y}]\to\mathbb{R}[\mathbf{x},\mathbf{y}]$ as follows:
$\operatorname{\mathbf{K}}_{2t}q(\mathbf{x},\mathbf{y})$ $:=\int_{\mathbf{X}}\operatorname{K}_{2t}((\mathbf{x},\mathbf{y}),(\mathbf{x}^{\prime},\mathbf{y}^{\prime});\lambda)q(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\operatorname{d}\mu_{\mathbf{X}}(\mathbf{x}^{\prime},\mathbf{y}^{\prime}),$ (16)
where $\text{where }\operatorname{K}_{2t}((\mathbf{x},\mathbf{y}),(\mathbf{x}^{\prime},\mathbf{y}^{\prime});\lambda)$ $:=\sum_{k,s=0}^{2t}\lambda_{k}\lambda_{s}\operatorname{K}^{(k,s)}((\mathbf{x},\mathbf{y}),(\mathbf{x}^{\prime},\mathbf{y}^{\prime})),$ (17)
and $\quad\text{and }\operatorname{K}^{(k,s)}((\mathbf{x},\mathbf{y}),(\mathbf{x}^{\prime},\mathbf{y}^{\prime}))$ $:=\sum_{|\alpha|=k,|\beta|=s}P_{\alpha}(\mathbf{x})P_{\beta}(\mathbf{y})P_{\alpha}(\mathbf{x}^{\prime})P_{\beta}(\mathbf{y}^{\prime}).$ (18)
Let us show that $\operatorname{\mathbf{K}}_{2t}$ satisfies the three desired properties (P1), (P2) and (P3).
For each $k,s$, one defines $H_{k,s}[\mathbf{x},\mathbf{y}]:=H_{k}[\mathbf{x}]\otimes H_{s}[\mathbf{y}]=\operatorname{span}\{P_{\alpha}(\mathbf{x})P_{\beta}(\mathbf{y}):\alpha,\beta\in\mathbb{N}^{n},|\alpha|=k,|\beta|=s\}$.
Then every $q\in\mathbb{R}[\mathbf{x},\mathbf{y}]_{d}$ has a decomposition
$q(\mathbf{x},\mathbf{y})=\sum_{k+s\leq d}q_{k,s}(\mathbf{x},\mathbf{y}),\text{ for all }(\mathbf{x},\mathbf{y})\in\mathbf{X},$ (19)
with $q_{k,s}\in H_{k,s}[\mathbf{x},\mathbf{y}]$, given by
$q_{k,s}=\int_{\mathbf{X}}\operatorname{C}^{(k)}(\mathbf{x},\mathbf{x}^{\prime})\operatorname{C}^{(s)}(\mathbf{y},\mathbf{y}^{\prime})q(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\operatorname{d}\mu(\mathbf{x}^{\prime}).$ (20)
By construction, one has
$\operatorname{K}^{(k,s)}((\mathbf{x},\mathbf{y}),(\mathbf{x}^{\prime},\mathbf{y}^{\prime}))$ $=\operatorname{C}^{(k)}(\mathbf{x},\mathbf{x}^{\prime})\operatorname{C}^{(s)}(\mathbf{y},\mathbf{y}^{\prime}),$
$\operatorname{K}_{2t}((\mathbf{x},\mathbf{y}),(\mathbf{x}^{\prime},\mathbf{y}^{\prime});\lambda)$ $=\sum_{k,s=0}^{2t}\lambda_{k}\lambda_{s}\operatorname{C}^{(k)}(\mathbf{x},\mathbf{x}^{\prime})\operatorname{C}^{(s)}(\mathbf{y},\mathbf{y}^{\prime})=\operatorname{C}_{2t}(\mathbf{x},\mathbf{x}^{\prime};\lambda)\operatorname{C}_{2t}(\mathbf{y},\mathbf{y}^{\prime};\lambda),$
where $\operatorname{C}_{2t}(\cdot,\cdot;\lambda)$ has been defined in (12). Then one has
$\operatorname{\mathbf{K}}_{2t}q=\sum_{k+s\leq d}\lambda_{k}\lambda_{s}q_{k,s}.$ (21)
Since $\lambda_{0}=1$, (P1) is satisfied.
To prove that (P2) is satisfied, recall that
$\operatorname{C}_{2t}(\mathbf{x},\mathbf{x}^{\prime};\lambda):=\sum_{k=0}^{2t}\lambda_{k}\operatorname{C}^{(k)}(\mathbf{x},\mathbf{x}^{\prime})=\sum_{k=0}^{2t}\lambda_{k}\mathcal{G}_{k}^{(n-1)}(\mathbf{x}\cdot\mathbf{x}^{\prime}),\quad\text{for all }\mathbf{x},\mathbf{x}^{\prime}\in S^{n-1},$
$\operatorname{C}_{2t}(\mathbf{y},\mathbf{y}^{\prime};\lambda):=\sum_{s=0}^{2t}\lambda_{s}\operatorname{C}^{(s)}(\mathbf{y},\mathbf{y}^{\prime})=\sum_{s=0}^{2t}\lambda_{s}\mathcal{G}_{s}^{(n-1)}(\mathbf{y}\cdot\mathbf{y}^{\prime}),\quad\text{for all }\mathbf{y},\mathbf{y}^{\prime}\in S^{n-1},$
and that we assumed that $\lambda=(\lambda_{k})$ is the set of coefficients of a univariate sum of squares in the basis of Gegenbauer polynomials.
Then at each fixed $(\mathbf{x}^{\prime},\mathbf{y}^{\prime})$, the polynomial $\operatorname{K}_{2t}((\mathbf{x},\mathbf{y}),(\mathbf{x}^{\prime},\mathbf{y}^{\prime});\lambda)$ is equal to a product of sums of squares of degree at most $4t$ on $\mathbf{X}$, so $\operatorname{K}_{2t}((\mathbf{x},\mathbf{y}),(\mathbf{x}^{\prime},\mathbf{y}^{\prime});\lambda)\in\mathcal{T}(\mathbf{X})_{4t}$ and Lemma 4 (when selecting $Q=\mathcal{T}(\mathbf{X})_{4t}$) implies that $\operatorname{\mathbf{K}}_{2t}$ satisfies (P2). Note that the degree $4t$ of this representation is twice larger than the one obtained in the case of the unit sphere.
We end the proof by showing that the operator satisfies (P3), namely,
$\max_{(\mathbf{x},\mathbf{y})\in\mathbf{X}}|\operatorname{\mathbf{K}}^{-1}q(\mathbf{x},\mathbf{y})-q(\mathbf{x},\mathbf{y})|\leq\varepsilon=\frac{C_{\mathbf{X}}(n,d)}{t^{2}},$ (22)
with
$C_{\mathbf{X}}(n,d):=8n^{2}d^{3}\binom{d+2}{2}\cdot\gamma(S^{n-1})_{d}^{2},$ (23)
and $\gamma(S^{n-1})_{d}$ is the constant defined in (13). By (21), one has
$\operatorname{\mathbf{K}}_{2t}^{-1}q-q=\sum_{1\leq k+s\leq d}\left(\frac{1}{\lambda_{k}\lambda_{s}}-1\right)q_{k,s}.$ (24)
To bound this quantity over $\mathbf{X}$, we follow a strategy similar to the one from [LS23], that relies on Bernoulli’s identity: for any $x\in[0,1]$ and $m\geq 1$, we have $1-(1-x)^{m}\leq mx$. For all $1\leq k\leq d$, let us define $\eta_{k}:=1-\lambda_{k}$, and $\eta:=\max_{1\leq k\leq d}\eta_{k}$. Since we selected $\lambda$ as in Lemma 6, one has $1/2\leq\lambda_{k}\leq 1$, for all $1\leq k\leq d$, then $20\leq\eta\leq 1/$ 2, so one can apply Bernoulli’s identity to obtain
$1-\lambda_{k}\lambda_{s}=1-(1-\eta_{k})(1-\eta_{s})\leq 1-(1-\eta^{2})\leq 2\eta.$ (25)
By (15), one has $1-\lambda_{k}\leq\sum_{k=1}^{d}(1-\lambda_{k})\leq\frac{n^{2}d^{3}}{t^{2}}$, for all $1\leq k\leq d$, implying that $\eta\leq\frac{n^{2}d^{3}}{t^{2}}$. Next, one has
$\left|1-\frac{1}{\lambda_{k}\lambda_{s}}\right|=\frac{|1-\lambda_{k}\lambda_{s}|}{|\lambda_{k}\lambda_{s}|}\leq 4(1-\lambda_{k}\lambda_{s}).$ (26)
Then by combining (25) and (26), we obtain
$\sum_{1\leq k+s\leq d}\left(\frac{1}{\lambda_{k}\lambda_{s}}-1\right)\leq 8\frac{n^{2}d^{3}}{t^{2}}\cdot\left|\mathbb{N}_{d}^{2}\right|=8\frac{n^{2}d^{3}}{t^{2}}\binom{d+2}{2}.$
Eventually let us define $p_{k}(\mathbf{x},\mathbf{y}):=\sum_{s}q_{k,s}(\mathbf{x},\mathbf{y})$. Let us fix $\mathbf{y}\in S^{n-1}$. From (19) one obtains the harmonic decomposition of $q(\mathbf{x},\mathbf{y})=\sum_{k}p_{k}(\mathbf{x},\mathbf{y})$, with $p_{k}\in H_{k}[\mathbf{x}]$. Therefore, one has
$\|p_{k}(\cdot,\mathbf{y})\|_{S^{n-1}}\leq\gamma(S^{n-1})_{d}\cdot\|q(\cdot,\mathbf{y})\|_{S^{n-1}}\leq\gamma(S^{n-1})_{d}.$
The above inequality is valid for all $\mathbf{y}\in S^{n-1}$, thus one has $\|p_{k}\|_{\mathbf{X}}\leq\gamma(S^{n-1})_{d}$. Now at each fixed $k$ and $\mathbf{x}$, one has $q_{k,s}(\mathbf{x},\cdot)\in H_{s}[\mathbf{y}]$, yielding the harmonic decomposition $p_{k}(\mathbf{x},\cdot)=\sum_{s}q_{k,s}(\mathbf{x},\cdot)$ and
$\|q_{k,s}(\mathbf{x},\cdot)\|_{S^{n-1}}\leq\gamma(S^{n-1})_{d}\cdot\|p_{k}(\mathbf{x},\cdot)\|_{S^{n-1}}\leq\gamma(S^{n-1})_{d}\cdot\|p_{k}\|_{\mathbf{X}}\leq\gamma(S^{n-1})^{2}_{d}.$
Since the above inequality is valid for all $\mathbf{x}\in S^{n-1}$, we obtain $\|q_{k,s}\|_{\mathbf{X}}\leq\gamma(S^{n-1})^{2}_{d}$, yielding the desired bound. ∎
Remark 7.
If one assumes that $q$ is bi-homogeneous polynomial of degree $d$, i.e., homogeneous of degree $d$ with respect to $\mathbf{x}$ (at fixed $\mathbf{y}$) and $\mathbf{y}$ (at fixed $\mathbf{x}$), then one can replace $\gamma(S^{n-1})^{2}_{d}$ by the quantity $\gamma(S^{n-1})^{2}_{=d}$ that can be bounded independently of the dimension $n$. This follows from [FF21, Proposition 5] which states that the supremum-norm of the harmonic components of a homogeneous polynomial can be bounded by a constant independent of the dimension.
3.2 Extension to arbitrary sphere products
One can easily extend the rate given in Theorem 3 when optimizing a polynomial over the Cartesian product of $m$ spheres $\mathbf{X}=S^{n-1}\underbrace{\times\dots\times}_{m-1}S^{n-1}$.
Theorem 8.
For $m\geq 2$, let $\mathbf{X}=S^{n-1}\times\dots\times S^{n-1}$ be the Cartesian product of $m$ unit spheres, and let $q\in\mathbb{R}[\mathbf{x}_{1},\dots,\mathbf{x}_{m}]_{d}$. Then for any $t\geq 2nd\sqrt{d}$, the lower bound $\operatorname{lb}(q,\mathcal{T}(\mathbf{X}))_{mt}$ for the minimization of $q$ over $\mathbf{X}$ satisfies:
$q_{\min}-\operatorname{lb}(q,\mathcal{T}(\mathbf{X}))_{mt}\leq\frac{C_{\mathbf{X}}(m,n,d)}{t^{2}}\cdot(q_{\max}-q_{\min}),$ (27)
with
$C_{\mathbf{X}}(n,d,m):=m2^{m}n^{2}d^{3}\binom{d+m}{m}\cdot\gamma(S^{n-1})_{d}^{m},$ (28)
and $\gamma(S^{n-1})_{d}$ is the constant defined in (13).
Proof.
Up to translation and scaling, we assume that $q_{\min}=0$ and $q_{\max}=1$. Let us fix $t\geq 2nd\sqrt{d}$. We outline how this rate is obtained by adapting the proof of Theorem 3 given in Section 3.1.
  • We equip $\mathbf{X}$ with the product measure $\mu_{\mathbf{X}}:=\mu\overbrace{\otimes\cdots\otimes}^{m-1}\mu$, where $\mu$ is the uniform Haar measure on $S^{n-1}$. The orthonormal basis with respect to $(\mathbf{X},\mu_{\mathbf{X}})$ is obtained by an $m$-th order tensorization;
  • For each $\kappa\in\mathbb{N}^{m}$, one defines $H_{\kappa}[\mathbf{x}_{1},\dots,\mathbf{x}_{m}]$ as the span of products of $m$ orthonormal polynomials that are homogeneous of degrees $\kappa_{1}$ in $\mathbf{x}_{1},\dots,\kappa_{m}$ in $\mathbf{x}_{m}$, respectively. Doing so, every polynomial $q\in\mathbb{R}[\mathbf{x}_{1},\dots,\mathbf{x}_{m}]_{d}$ has a decomposition $q=\sum_{|\kappa|\leq d}q_{\kappa}$, with $q_{\kappa}\in H_{\kappa}[\mathbf{x}_{1},\dots,\mathbf{x}_{m}]$;
  • The perturbed kernel $\operatorname{K}_{2t}(\cdot,\cdot;\lambda)$ is obtained by multiplying $m$ kernels $\operatorname{C}_{2t}(\cdot,\cdot;\lambda):S^{n-1}\times S^{n-1}\to\mathbb{R}$, with $\lambda=(\lambda_{k})$ as in Lemma 6. The polynomial obtained at each fixed second argument belongs to $\mathcal{T}(\mathbf{X})_{2mt}$ and Lemma 4 implies that the operator $\operatorname{\mathbf{K}}_{2t}$ satisfies (P2);
  • Let us define $\lambda_{\kappa}:=\lambda_{\kappa_{1}}\cdots\lambda_{\kappa_{m}}$. Since $\lambda_{0}=1$, (P1) is satisfied. In addition, one obtains
    $\operatorname{\mathbf{K}}_{2t}^{-1}q-q=\sum_{1\leq|\kappa|\leq d}\left(\frac{1}{\lambda_{\kappa}}-1\right)q_{\kappa}.$ (29)
    As for (25), we rely on Bernoulli’s identity to obtain the inequality $1-\lambda_{\kappa}\leq m\eta$, with $\eta\leq\frac{n^{2}d^{3}}{t^{2}}$. Since $\lambda_{k}\geq 1/2$, one has $|1-\lambda_{\kappa}^{-1}|\leq 2^{m}(1-\lambda_{\kappa})$. Eventually, one has $\|q_{\kappa}\|_{\mathbf{X}}\leq\gamma(S^{n-1})^{m}_{d}$.
3.3 Extension to products of distinct sets
Given a compact set $\mathbf{X}\subset\mathbb{R}^{n}$, we denote by $\mathcal{M}(\mathbf{X})$ the space of positive Borel measures supported on $\mathbf{X}$. In this section, we assume that convergence rates have been obtained for polynomial optimization over two given sets $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$, thanks to perturbed kernels and associated operators satisfying (P1), (P2) and (P3). We explain how to derive a convergence rate for polynomial optimization over the Cartesian product $\mathbf{X}=\mathbf{X}_{1}\times\mathbf{X}_{2}$.
Theorem 9.
Let $\mathbf{X}_{1},\mathbf{X}_{2}\subseteq\mathbb{R}^{n}$ be two compact semialgebraic sets, and let $\mathbf{X}=\mathbf{X}_{1}\times\mathbf{X}_{2}$ be their Cartesian product. For each $i=1,2$, assume that there exist
  • $m_{i}\in\mathbb{N}$, $t_{i}\in\mathbb{R}_{\geq 0}$,
  • a function $d_{i}:\mathbb{N}\to\mathbb{N}$,
  • a probability measure $\mu_{i}\in\mathcal{M}(\mathbf{X}_{i})$,
  • a function $\eta^{(i)}:\mathbb{N}\to\mathbb{R}_{\geq 0}$, converging to 00 at infinity,
  • a decomposition $\mathcal{P}(\mathbf{X}_{i})=\displaystyle\bigoplus_{\kappa^{(i)}\in\mathbb{N}^{m_{i}}}H_{\kappa^{(i)}}$, where each $H_{\kappa^{(i)}}$ spans a subspace $B_{\kappa^{(i)}}$ of orthonormal polynomials with respect to $(\mathbf{X}_{i},\mu_{i})$,
  • a vector $\lambda^{(i)}=\left(\lambda_{\kappa^{(i)}}^{(i)}\right)$ indexed by $\mathbb{N}_{2t}^{m_{i}}$, for all $t\geq t_{i}$,
such that the following hold
  1. 1.
    the polynomial kernel $\operatorname{C}_{i,2t}(\cdot,\cdot;\lambda^{(i)}):\mathbf{X}_{i}\times\mathbf{X}_{i}\to\mathbb{R}$ defined for all $t\geq t_{i}$ by
    $\operatorname{C}_{i,2t}(\mathbf{x},\mathbf{x}^{\prime};\lambda^{(i)}):=\sum_{|\kappa^{(i)}|\leq 2t}\lambda_{\kappa^{(i)}}^{(i)}\sum_{P\in B_{\kappa^{(i)}}}P(\mathbf{x})P(\mathbf{x}^{\prime}),$
    satisfies $\operatorname{C}_{i,2t}(\cdot,\mathbf{x}^{\prime};\lambda^{(i)})\in\mathcal{T}(\mathbf{X}_{i})_{d_{i}(t)}$, at each fixed $\mathbf{x}^{\prime}\in\mathbf{X}_{i}$, for all $t\geq t_{i}$,
  2. 2.
    the eigenvalues of the linear operator $\operatorname{\mathbf{C}}_{i,2t}:\mathbb{R}[\mathbf{x}_{i}]\to\mathbb{R}[\mathbf{x}_{i}]$ defined for all $t\geq t_{i}$ by
    $\operatorname{\mathbf{C}}_{i,2t}p:=\displaystyle\int_{\mathbf{X}_{i}}\operatorname{C}_{i,2t}(\mathbf{x},\mathbf{x}^{\prime})p(\mathbf{x}^{\prime})\operatorname{d}\mu_{i}(\mathbf{x}^{\prime}),$
    satisfy
    $\lambda_{0}^{(i)}=1,\quad 1/2\leq\lambda_{\kappa^{(i)}}^{(i)}\leq 1\quad\text{for all }\kappa^{(i)}\in\mathbb{N}_{2t}^{m_{i}}\text{ and }t\geq t_{i},$ (30)
    $1-\lambda_{\kappa^{(i)}}^{(i)}\leq\eta^{(i)}(t)\quad\text{for all }t\geq t_{i},\text{ and for all }1\leq|\kappa^{(i)}|\leq 2t.$ (31)
Define $m:=m_{1}+m_{2}$ and $\eta(t):=\max\{\eta^{(1)}(t),\eta^{(2)}(t)\}$. Let $q\in\mathbb{R}[\mathbf{x}_{1},\mathbf{x}_{2}]_{d}$. Then for any $t\geq\max\{t_{1},t_{2}\}$, the lower bound $\operatorname{lb}(q,\mathcal{T}(\mathbf{X}))_{mt}$ for the minimization of $q$ over $\mathbf{X}$ satisfies:
$q_{\min}-\operatorname{lb}(q,\mathcal{T}(\mathbf{X}))_{d_{1}(t)+d_{2}(t)}\leq 8\eta(t)\cdot\binom{m+d}{d}\cdot\gamma(\mathbf{X}_{1})_{d}\cdot\gamma(\mathbf{X}_{2})_{d}\cdot(q_{\max}-q_{\min}),$ (32)
Proof.
Up to translation and scaling, we assume that $q_{\min}=0$ and $q_{\max}=1$. Let us fix $t\geq\max\{t_{1},t_{2}\}$. We outline how this rate is obtained by adapting the proof of Theorem 3 given in Section 3.1.
  • We equip $\mathbf{X}$ with the product measure $\mu_{\mathbf{X}}:=\mu_{1}\otimes\mu_{2}$;
  • For each $\kappa=(\kappa^{(1)},\kappa^{(2)})\in\mathbb{N}^{m}$, one defines $H_{\kappa}[\mathbf{x}_{1},\mathbf{x}_{2}]:=H_{\kappa^{(1)}}[\mathbf{x}_{1}]\otimes H_{\kappa^{(2)}}[\mathbf{x}_{2}]$. Doing so, every polynomial $q\in\mathbb{R}[\mathbf{x}_{1},\mathbf{x}_{2}]_{d}$ has a decomposition $q=\sum_{|\kappa|\leq d}q_{\kappa}$, with $q_{\kappa}\in H_{\kappa}[\mathbf{x}_{1},\mathbf{x}_{2}]$;
  • For each $\kappa=(\kappa^{(1)},\kappa^{(2)})\in\mathbb{N}^{m}$, one defines $\lambda_{\kappa}:=\lambda_{\kappa^{(1)}}\lambda_{\kappa^{(2)}}$. We build a linear invertible operator $\operatorname{\mathbf{K}}_{2t}:\mathbb{R}[\mathbf{x}_{1},\mathbf{x}_{2}]\to\mathbb{R}[\mathbf{x}_{1},\mathbf{x}_{2}]$ as follows:
    $\operatorname{\mathbf{K}}_{2t}q(\mathbf{x}_{1},\mathbf{x}_{2})$ $:=\int_{\mathbf{X}}\operatorname{K}_{2t}((\mathbf{x}_{1},\mathbf{x}_{2}),(\mathbf{x}_{1}^{\prime},\mathbf{x}_{2}^{\prime});\lambda)q(\mathbf{x}_{1}^{\prime},\mathbf{x}_{2}^{\prime})\operatorname{d}\mu_{\mathbf{X}}(\mathbf{x}_{1}^{\prime},\mathbf{x}_{2}^{\prime}),$
    where $\text{where }\operatorname{K}_{2t}((\mathbf{x}_{1},\mathbf{x}_{2}),(\mathbf{x}_{1}^{\prime},\mathbf{x}_{2}^{\prime});\lambda)$ $:=\operatorname{C}_{1,2t}(\mathbf{x}_{1},\mathbf{x}_{1}^{\prime};\lambda^{(1)})\operatorname{C}_{2,2t}(\mathbf{x}_{2},\mathbf{x}_{2}^{\prime};\lambda^{(2)}).$
    Then, at each fixed $(\mathbf{x}_{1}^{\prime},\mathbf{x}_{2}^{\prime})\in\mathbf{X}$, $\operatorname{C}_{2t}(\cdot,(\mathbf{x}_{1}^{\prime},\mathbf{x}_{2}^{\prime});\lambda)\in\mathcal{T}(\mathbf{X})_{d_{1}(t)+d_{2}(t)}$ and $\operatorname{\mathbf{K}}_{2t}$ satisfies (P2);
  • Since $\lambda_{0}=1$, $\operatorname{\mathbf{K}}_{2t}$ satisfies (P1);
  • One has
    $\operatorname{\mathbf{K}}_{2t}^{-1}q-q=\sum_{|\kappa|\leq d}\left(\frac{1}{\lambda_{\kappa}}-1\right)q_{\kappa}.$ (33)
    As for (25), we rely on Bernoulli’s identity to obtain the inequality $1-\lambda_{\kappa}\leq 2\eta(t)$. Since $\lambda_{k}\geq 1/2$, one has $|1-\lambda_{\kappa}^{-1}|\leq 4(1-\lambda_{\kappa})$. Eventually, one has $\|q_{\kappa}\|_{\mathbf{X}}\leq\gamma(\mathbf{X}_{1})_{d}\cdot\gamma(\mathbf{X}_{2})_{d}$, so $\operatorname{\mathbf{K}}_{2t}$ satisfies (P3).
We now give examples of sets where the assumptions of Theorem 9 are fulfilled.
The unit sphere [FF21], the unit ball [Slo22] or the standard simplex[Slo22]
[…]
  • $m_{i}=1$, $t_{i}=2nd\sqrt{d}$;
  • $d_{i}:=t\mapsto 2t$;
  • $\mu_{i}$ to be either the uniform Haar measure for $S^{n-1}$, or the measure with density $c_{n}(1-\|\mathbf{x}\|^{2})^{-\frac{1}{2}}$ (with $c_{n}>0$ being a normalization constant) for $\mathbb{B}^{n}$, or the measure with density $c_{n}x_{1}^{-1/2}\cdots x_{n}^{-1/2}(1-\sum_{i=1}^{n}x_{i})^{-\frac{1}{2}}$ (with $c_{n}>0$ being a normalization constant) for $\Delta^{n}$;
  • $\eta^{(i)}:t\mapsto\frac{n^{2}d^{3}}{t^{2}}$;
  • $\mathcal{P}(S^{n-1})=\displaystyle\bigoplus_{\kappa^{(i)}=0}^{\infty}H_{\kappa^{(i)}}$, where $H_{\kappa^{(i)}}$ is the spanning set of orthonormal polynomials with respect to $(S^{n-1},\mu_{i})$ of exact degree $\kappa^{(i)}$;
  • the vector $\lambda^{(i)}=\left(\lambda_{\kappa^{(i)}}^{(i)}\right)$, indexed by $\mathbb{N}_{2t}$, for all $t\geq t_{i}$, as in Lemma 6.
The hypercube [LS23]
If $\mathbf{X}_{i}=[-1,1]^{n}$, then one can select
  • $m_{i}=n$, $t_{i}=\pi d\sqrt{2n}$;
  • $d_{i}:t\mapsto n(t+1)$;
  • $\mu_{i}$ to be the normalized Chebyshev measure with density $\frac{1}{\pi\sqrt{1-x_{1}^{2}}}\cdots\frac{1}{\pi\sqrt{1-x_{n}^{2}}}$;
  • $\eta^{(i)}:t\mapsto\frac{n\pi^{2}d^{2}}{t^{2}}$,
  • $\mathcal{P}([-1,1]^{n})=\displaystyle\bigoplus_{\kappa^{(i)}\in\mathbb{N}^{n}}H_{\kappa^{(i)}}$, where $H_{\kappa^{(i)}}=\mathbb{R}T_{\kappa^{(i)}}$ and $T_{\kappa^{(i)}}$ is the multivariate Chebyshev polynomial defined as the product of univariate Chebyshev polynomials:
    $T_{\kappa^{(i)}}(\mathbf{x}):=\prod_{j=1}^{n}T_{\kappa_{j}^{(i)}}(x_{j}),\quad T_{k}(\cos\phi):=\cos(k\phi)\quad(\phi\in\mathbb{R},k\in\mathbb{N});$
  • the vector $\lambda^{(i)}=\left(\lambda_{\kappa^{(i)}}^{(i)}\right)$, indexed by $\mathbb{N}^{n}_{2t}$, for all $t\geq t_{i}$, as in [LS23, Lemma 12].
4 Application to quantum Wasserstein distance
In this section we analyze the convergence rate of the hierarchy derived in [CM25] to approximate as closely as desired the so-called order 2 quantum Wasserstein distance. Roughly speaking, the classical Wasserstein distance quantifies the cost of transporting one probability measure to another. In order to extend this notion to quantum states, several notions of quantum Wasserstein distances have been proposed; see, e.g., [Nie06] and [DPMTL21]. In [CM25], the authors specifically focus on the order 2 quantum Wasserstein distance that has been recently proposed in [BF24].
4.1 Problem statement and main result
Additional complex notation.
Let $\mathbf{i}=\sqrt{-1}\in\mathbb{C}$. Given $x\in\mathbb{C}$, one can write $x=a+\mathbf{i}\,b$, where $a=\operatorname{Re}(x)$ is the real part of $x$ and $b=\operatorname{Im}(x)$ is the imaginary part of $x$. The complex conjugate of $x$ is $\overline{x}=a-\mathbf{i}\,b$. We extend this to vectors and matrices by applying the corresponding operations entrywise. The modulus of $x$ is denoted by $|x|:=\sqrt{a^{2}+b^{2}}$. The set of polynomials in variables $\mathbf{x}=(x_{1},\dots,x_{n})$ and $\overline{\mathbf{x}}=(\overline{x}_{1},\dots,\overline{x}_{n})$ with complex coefficients is denoted by $\mathbb{C}[\mathbf{x},\overline{\mathbf{x}}]$. A polynomial $p\in\mathbb{C}[\mathbf{x},\overline{\mathbf{x}}]$ satisfying $p=\overline{p}$ is called Hermitian. Given a vector $u\in\mathbb{C}^{n}$, $u^{*}$ is the row vector containing the conjugate entries of $u$. We will use the same notation as for the real case for the Euclidean norm $\|u\|=\sqrt{u^{*}u}$. The $\ell_{1}$-norm $\|u\|_{1}$ is the sum of moduli of the entries of $u$. A linear functional $L:\mathbb{C}[\mathbf{x},\overline{\mathbf{x}}]\to\mathbb{C}$ is called Hermitian if $\overline{L(p)}=L(\overline{p})$ for all $p\in\mathbb{C}[\mathbf{x},\overline{\mathbf{x}}]$. Given a compact set $\mathbf{X}\subset\mathbb{C}^{n}$, we denote by $\mathcal{M}(\mathbf{X})$ the space of complex positive Borel measures supported on $\mathbf{X}$.
Given $n\in\mathbb{N}$, let $\rho,\nu\in\mathbb{C}^{n\times n}$ be two normalized quantum states, i.e., Hermitian matrices with nonnegative eigenvalues and unit trace. A quantum transport plan between $\rho$ and $\nu$ is a finite set of triples $\{(\omega_{\ell},u_{\ell},v_{\ell})\}_{\ell}$ such that
$\sum_{\ell}\omega_{\ell}\,u_{\ell}u_{\ell}^{*}=\rho\quad\text{ and}\quad\sum_{j}\omega_{\ell}\,v_{\ell}v_{\ell}^{*}=\nu,$ (34)
where $\omega_{\ell}>0$, $\sum_{\ell}\omega_{\ell}=1$, $u_{\ell},v_{\ell}\in\mathbb{C}^{n}$, $\|u_{\ell}\|=\|v_{\ell}\|=1$. The set of all such quantum transport plans between two states $\rho$ and $\nu$ is denoted by $\mathcal{Q}(\rho,\nu)$. Given a plan $e=\{(\omega_{\ell},u_{\ell},v_{\ell})\}_{\ell}$, the associated (order $2$) quantum transport cost is defined as
$T_{2}(e):=\sum_{\ell}\omega_{\ell}\operatorname{Tr}\left[(u_{\ell}u_{\ell}^{*}-v_{\ell}v_{\ell}^{*})\overline{(u_{\ell}u_{\ell}^{*}-v_{\ell}v_{\ell}^{*})}\right],$ (35)
and the order $2$ quantum Wasserstein distance is defined as
$W_{2}(\rho,\nu):=\Big{(}\inf_{e\in\mathcal{Q}(\rho,\nu)}T_{2}(e)\Big{)}^{1/2}.$ (36)
In [GLS22, Section 5.2], the authors propose a moment reformulation of the DPS hierarchy [DPS02], initially designed to distinguish separable and entangled states. Inspired by this reformulation, it is proved in [CM25, Theorem 8] that $W_{2}^{2}(\rho,\nu)$ is the optimal value of a particular instance of the generalized moment problem (GMP), i.e., an infinite-dimensional linear problem over positive Borel measures:
$W^{2}_{2}(\rho,\sigma)=\inf_{\mu\in\mathcal{M}(\mathbf{X})}$ $\int_{\mathbf{X}}f\operatorname{d}\mu$ (37)
$\operatorname{s.t.}$ $\int_{\mathbf{X}}\mathbf{x}\mathbf{x}^{*}\operatorname{d}\mu=\rho,$
$\int_{\mathbf{X}}\mathbf{y}\mathbf{y}^{*}\operatorname{d}\mu=\nu,$
where $f(\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}})=\operatorname{Tr}[(\mathbf{x}\mathbf{x}^{*}-\mathbf{y}\mathbf{y}^{*})\overline{(\mathbf{x}\mathbf{x}^{*}-\mathbf{y}\mathbf{y}^{*})}]$ and $\mathbf{X}=\{(\mathbf{x},\mathbf{y})\in\mathbb{C}^{2n}:\|\mathbf{x}\|=\|\mathbf{y}\|=1\}$ is the Cartesian product of two complex unit spheres. The dual of problem (37) is given by
$\sup_{\Lambda,\Gamma\in\mathbb{C}^{n\times n}}$ $\operatorname{Tr}(\rho\Lambda+\nu\Gamma)$ (38)
s.t. $f-\mathbf{x}^{*}\Lambda\mathbf{x}-\mathbf{y}^{*}\Gamma\mathbf{y}\geq 0\text{ on }\mathbf{X},$
$\Lambda^{*}=\Lambda,\quad\Gamma^{*}=\Gamma.$
The Schmüdgen-type moment hierarchy for (37) is given by
$W^{2}_{2}(\rho,\sigma)_{t}:=\inf_{L}$ $L(f)$ (39)
s.t. $L:\mathbb{C}[\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}}]_{2t}\to\mathbb{C}\text{ Hermitian linear},$
$L(\mathbf{x}\mathbf{x}^{*})=\rho,$
$L(\mathbf{y}\mathbf{y}^{*})=\nu,$
$L\geq 0\text{ on }\mathcal{T}(\mathbf{X})_{2t}.$
It is proved in [CM25, Theorem 10] that the corresponding sequence of lower bounds given by this hierarchy converges to the order $2$ quantum Wasserstein distance, i.e., $W^{2}_{2}(\rho,\sigma)_{t}\to W^{2}_{2}(\rho,\sigma)$ when $t$ goes to infinity. In the recent contribution [GM25], the authors analyze convergence rates of hierarchies of lower bounds for the GMP. We recall this result in the case where the GMP instance involves a single measure and equality constraints only.
Lemma 10 (See Theorem 3.4 and Corollary 3.5 from [GM25]).
Given $n,d,N\in\mathbb{N}$, $\tau\in\mathbb{R}^{N}$, $f,h_{1},\dots,h_{N}\in\mathbb{R}[\mathbf{x}]_{d}$, a compact semialgebraic set $\mathbf{X}\subset\mathbb{R}^{n}$, let us consider the GMP instance
$\mathfrak{p}:=\inf_{\mu\in\mathcal{M}(\mathbf{X})}$ $\int_{\mathbf{X}}f\operatorname{d}\mu$ (40)
$\operatorname{s.t.}$ $\int_{\mathbf{X}}h_{i}\operatorname{d}\mu=\tau_{i},\quad i=1,\dots,N,$
assumed to have a nonempty feasible set. The associated moment hierarchy of Schmüdgen-type bounds is given by
$\mathfrak{p}_{t}:=\inf_{L}$ $L(f)$ (41)
$\operatorname{s.t.}$ $L:\mathbb{R}[\mathbf{x}]_{2t}\to\mathbb{R}\text{ linear},$
$L(h_{i})=\tau_{i},\quad i=1,\dots,N,$
$L\geq 0\text{ on }\mathcal{T}(\mathbf{X})_{2t}.$
Assume that $\mathcal{T}(\mathbf{X})$ is Archimedean, and that a convergence rate is given for the hierarchy of Schmüdgen-type bounds for polynomial optimization on $\mathbf{X}$, i.e., there exists nonnegative constants (depending on $n$, $d$, $\mathbf{X}$) $C_{\mathbf{X}}$ and $t_{0}$ such that for all $q\in\mathbb{R}[\mathbf{x}]_{d}$ and $t\geq t_{0}$, one has $q_{\min}-\operatorname{lb}(q,\mathcal{T}(\mathbf{X}))_{t}\leq C_{\mathbf{X}}\eta(t)\cdot(q_{\max}-q_{\min})$, with $\eta:\mathbb{N}\to\mathbb{R}_{\geq 0}$ being a function converging to 00 at infinity. Assume that there exists $w\in\mathbb{R}^{n}$ such that $\sum_{i=1}^{N}w_{i}h_{i}>0$ on $\mathbf{X}$.
Then for all $t\geq t_{0}$, one has
$0\leq\mathfrak{p}-\mathfrak{p}_{t}\leq\frac{1+\sum_{i=1}^{N}\tau_{i}w_{i}}{(\sum_{i=1}^{N}w_{i}h_{i})_{\min}}C_{\mathbf{X}}\eta(t)\left(3\|f\|_{\mathbf{X}}+\|w^{\operatorname{opt}}\|_{1}\|h\|_{\mathbf{X}}\right),$
where $\|h\|_{\mathbf{X}}:=\max_{1\leq i\leq N}\|h_{i}\|_{\mathbf{X}}$, and $w^{\operatorname{opt}}$ is a vector achieving the supremum in the dual of (40).
Our final result, given in Theorem 11 below, provides the convergence rate of the moment hierarchy (39) for the order $2$ quantum Wasserstein distance, assuming that there is a bounded dual optimal solution of (38). This result is obtained by combining Lemma 10 with our analysis derived in Theorem 3.
Theorem 11.
Let assume that the dual (38) is attained at a bounded tuple. For any $t\geq 32n$, one has
$0\leq W^{2}_{2}(\rho,\nu)-W^{2}_{2}(\rho,\nu)_{2t}\leq\frac{\kappa(n,\rho,\nu)}{t^{2}},$ (42)
where $\kappa(n,\rho,\nu)$ is a constant depending only on $n$, $\rho$ and $\nu$, and being proportional to the constant $C_{\mathbf{X}}(2n,4)$ from Theorem 3. See relation (50) for details.
The proof of Theorem 11 is postponed in Section 4.3.
4.2 Changing variables from complex to real
In order to apply Lemma 10, we provide an equivalent formulation of the moment hierarchy (39) involving linear functionals acting on real polynomials. This reformulation is obtained thanks to [GLS22, Appendix A], where the authors describe how to transform polynomials in $\mathbb{C}[\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}}]$ into real polynomials in $\mathbb{R}[\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}]$ via the change of variables $\mathbf{x}=\mathbf{a}+\mathbf{i}\,\mathbf{b}$, $\mathbf{y}=\mathbf{c}+\mathbf{i}\,\mathbf{d}$. In this way, any polynomial $p\in\mathbb{C}[\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}}]$ corresponds to a unique pair of real polynomials
$p_{\operatorname{Re}}(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d})=\operatorname{Re}(p(\mathbf{a}+\mathbf{i}\,\mathbf{b},\mathbf{a}-\mathbf{i}\,\mathbf{b},\mathbf{c}+\mathbf{i}\,\mathbf{d},\mathbf{c}-\mathbf{i}\,\mathbf{d})),$ (43)
$p_{\operatorname{Im}}(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d})=\operatorname{Im}(p(\mathbf{a}+\mathbf{i}\mathbf{b},\mathbf{a}-\mathbf{i}\,\mathbf{b},\mathbf{c}+\mathbf{i}\,\mathbf{d},\mathbf{c}-\mathbf{i}\,\mathbf{d})),$ (44)
satisfying $p(\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}})=p_{\operatorname{Re}}(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d})+\mathbf{i}\,p_{\operatorname{Im}}(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d})$. When $p$ is Hermitian, one has $p(\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}})=p_{\operatorname{Re}}(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d})$, so in particular one has
$f(\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}})=\operatorname{Tr}[(\mathbf{x}\mathbf{x}^{*}-\mathbf{y}\mathbf{y}^{*})\overline{(\mathbf{x}\mathbf{x}^{*}-\mathbf{y}\mathbf{y}^{*})}]=f_{\operatorname{Re}}(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}).$
Given a set of Hermitian polynomials, we define its real analog by applying the map $\operatorname{Re}(\cdot)$ entrywise to this set. In particular, the real analog of the complex bi-sphere $\mathbf{X}$ is the set $\mathbf{X}_{\operatorname{Re}}=\{(\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d})\in\mathbb{R}^{4n}:1-\|\mathbf{a}\|^{2}-\|\mathbf{b}\|^{2}=1-\|\mathbf{c}\|^{2}-\|\mathbf{d}\|^{2}=0\}$. The associated truncated real quadratic module is
$\mathcal{T}^{\mathbb{R}}(\mathbf{X}_{\operatorname{Re}})_{2t}:=\left\{\sigma+q_{1}(1-\|\mathbf{a}\|^{2}-\|\mathbf{b}\|^{2})+q_{2}(1-\|\mathbf{c}\|^{2}-\|\mathbf{d}\|^{2}):\right.$ $\sigma\in\Sigma[\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}]_{2t},$
$\left.q_{1},q_{2}\in\mathbb{R}[\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}]_{2t-2}\right\}.$
Similarly, the maps $\operatorname{Re}(\cdot)$ and $\operatorname{Im}(\cdot)$ act entrywise for vectors and matrices with polynomial entries in $\mathbb{C}[\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}}]$, e.g., $\operatorname{Re}(\mathbf{x}\mathbf{x}^{*})=\mathbf{a}\mathbf{a}^{T}+\mathbf{b}\mathbf{b}^{T}$ and $\operatorname{Im}(\mathbf{x}\mathbf{x}^{*})=\mathbf{b}\mathbf{a}^{T}-\mathbf{a}\mathbf{b}^{T}$. Eventually, for a linear functional $L:\mathbb{C}[\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}}]\to\mathbb{C}$, one has $L(p)=\operatorname{Re}(L(p))+\mathbf{i}\,\operatorname{Im}(L(p))$. If in addition $L$ is Hermitian, that is, satisfies $\overline{L(p)}=L(\overline{p})$, then one can associate to $L$ a real linear functional $L^{\mathbb{R}}:\mathbb{R}[\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}]\to\mathbb{R}$, defined by
$L^{\mathbb{R}}(q)=L\left(q\left(\frac{\mathbf{x}+\overline{\mathbf{x}}}{2},\frac{\mathbf{x}-\overline{\mathbf{x}}}{2\mathbf{i}},\frac{\mathbf{y}+\overline{\mathbf{y}}}{2},\frac{\mathbf{y}-\overline{\mathbf{y}}}{2\mathbf{i}}\right)\right).$
Then for any $p\in\mathbb{C}[\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}}]$, one has $L(p)=L^{\mathbb{R}}(p_{\operatorname{Re}})+\mathbf{i}\,L^{\mathbb{R}}(p_{\operatorname{Im}})$. For any $t\geq 2$, we can now provide the following relaxation over real linear functionals:
$\inf_{L^{\mathbb{R}}}$ $L^{\mathbb{R}}(f^{\operatorname{Re}})$ (45)
s.t. $L^{\mathbb{R}}:\mathbb{R}[\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}]_{2t}\to\mathbb{R}\text{ linear},$
$L^{\mathbb{R}}(\mathbf{a}\mathbf{a}^{T}+\mathbf{b}\mathbf{b}^{T})=\operatorname{Re}(\rho),$
$L^{\mathbb{R}}(\mathbf{b}\mathbf{a}^{T}-\mathbf{a}\mathbf{b}^{T})=\operatorname{Im}(\rho),$
$L^{\mathbb{R}}(\mathbf{c}\mathbf{c}^{T}+\mathbf{d}\mathbf{d}^{T})=\operatorname{Re}(\nu),$
$L^{\mathbb{R}}(\mathbf{d}\mathbf{c}^{T}-\mathbf{c}\mathbf{d}^{T})=\operatorname{Im}(\nu),$
$L^{\mathbb{R}}\geq 0\text{ on }\mathcal{T}^{\mathbb{R}}(\mathbf{X}_{\operatorname{Re}})_{2t}.$
As a consequence of the above defined transformations, the relaxation (45) is equivalent to the moment relaxation (39) over Hermitian linear functionals. In particular the equivalence
$L\geq 0\text{ on }\mathcal{T}(\mathbf{X})_{2t}\iff L^{\mathbb{R}}\geq 0\text{ on }\mathcal{T}^{\mathbb{R}}(\mathbf{X}_{\operatorname{Re}})_{2t}$
is proved in [GLS22, Lemma 33]. As in the complex setting, one can show that the hierarchy of relaxations converges to $W^{2}_{2}(\rho,\nu)$ when $t$ goes to infinity, that is
$W^{2}_{2}(\rho,\nu)=\inf_{L^{\mathbb{R}}}$ $L^{\mathbb{R}}(f_{\operatorname{Re}})$ (46)
s.t. $L^{\mathbb{R}}:\mathbb{R}[\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}]\to\mathbb{R}\text{ linear},$
$L^{\mathbb{R}}(\mathbf{a}\mathbf{a}^{T}+\mathbf{b}\mathbf{b}^{T})=\operatorname{Re}(\rho),$
$L^{\mathbb{R}}(\mathbf{b}\mathbf{a}^{T}-\mathbf{a}\mathbf{b}^{T})=\operatorname{Im}(\rho),$
$L^{\mathbb{R}}(\mathbf{c}\mathbf{c}^{T}+\mathbf{d}\mathbf{d}^{T})=\operatorname{Re}(\nu),$
$L^{\mathbb{R}}(\mathbf{d}\mathbf{c}^{T}-\mathbf{c}\mathbf{d}^{T})=\operatorname{Im}(\nu),$
$L^{\mathbb{R}}\geq 0\text{ on }\mathcal{T}^{\mathbb{R}}(\mathbf{X}_{\operatorname{Re}}).$
The latter formulation can be written equivalently as an infinite-dimensional linear problem over real positive Borel measures supported on $\mathbf{X}_{\operatorname{Re}}$:
$W^{2}_{2}(\rho,\nu)=\inf_{\mu^{\mathbb{R}}\in\mathcal{M}(\mathbf{X}_{\operatorname{Re}})}$ $\int f_{\operatorname{Re}}\operatorname{d}\mu^{\mathbb{R}}$ (47)
s.t. $\int(\mathbf{a}\mathbf{a}^{T}+\mathbf{b}\mathbf{b}^{T})\operatorname{d}\mu^{\mathbb{R}}=\operatorname{Re}(\rho),$
$\int(\mathbf{b}\mathbf{a}^{T}-\mathbf{a}\mathbf{b}^{T})\operatorname{d}\mu^{\mathbb{R}}=\operatorname{Im}(\rho),$
$\int(\mathbf{c}\mathbf{c}^{T}+\mathbf{d}\mathbf{d}^{T})\operatorname{d}\mu^{\mathbb{R}}=\operatorname{Re}(\nu),$
$\int(\mathbf{d}\mathbf{c}^{T}-\mathbf{c}\mathbf{d}^{T})\operatorname{d}\mu^{\mathbb{R}}=\operatorname{Im}(\nu).$
Note that every feasible solution $\mu_{\mathbb{R}}$ of (47) must be a probability measure. Indeed, combining the first equality constraint together with the fact that $\mu_{\mathbb{R}}$ is supported on $\mathbf{X}_{\operatorname{Re}}$ yields
$\int\operatorname{Tr}(\mathbf{a}\mathbf{a}^{T}+\mathbf{b}\mathbf{b}^{T})\operatorname{d}\mu^{\mathbb{R}}=\int(\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2})\operatorname{d}\mu^{\mathbb{R}}=\int\operatorname{d}\mu^{\mathbb{R}}=\operatorname{Tr}(\operatorname{Re}(\rho))=1.$
The dual of (47) is given by
$\sup_{\Lambda_{i},\Gamma_{i}\in\mathbb{R}^{n\times n}}$ $\operatorname{Tr}(\operatorname{Re}(\rho)\Lambda_{1}+\operatorname{Re}(\nu)\Gamma_{1}+\operatorname{Im}(\rho)\Lambda_{2}+\operatorname{Im}(\nu)\Gamma_{2})$ (48)
s.t. $f_{\operatorname{Re}}\geq\operatorname{Tr}((\mathbf{a}\mathbf{a}^{T}+\mathbf{b}\mathbf{b}^{T})\Lambda_{1}+(\mathbf{c}\mathbf{c}^{T}+\mathbf{d}\mathbf{d}^{T})\Gamma_{1}+(\mathbf{b}\mathbf{a}^{T}-\mathbf{a}\mathbf{b}^{T})\Lambda_{2}$
$\quad\quad\quad+(\mathbf{d}\mathbf{c}^{T}-\mathbf{c}\mathbf{d}^{T})\Gamma_{2})\text{ on }\mathbf{X}_{\operatorname{Re}},$
$\Lambda_{1}^{T}=\Lambda_{1},\quad\Lambda_{2}^{T}=-\Lambda_{2},\quad\Gamma_{1}^{T}=\Gamma_{1},\quad\Gamma_{2}^{T}=-\Gamma_{2}.$
4.3 Proof of Theorem 11
We are now in the appropriate setup to apply Lemma 10.
Proof.
The three required assumptions are as follows:
  1. 1.
    the instance of the generalized moment problem has a non-empty feasible set;
  2. 2.
    there exists $(\Lambda_{i},\Gamma_{i})$ such that $\operatorname{Tr}((\mathbf{a}\mathbf{a}^{T}+\mathbf{b}\mathbf{b}^{T})\Lambda_{1}+(\mathbf{c}\mathbf{c}^{T}+\mathbf{d}\mathbf{d}^{T})\Gamma_{1}+(\mathbf{b}\mathbf{a}^{T}-\mathbf{a}\mathbf{b}^{T})\Lambda_{2}+(\mathbf{d}\mathbf{c}^{T}-\mathbf{c}\mathbf{d}^{T})\Gamma_{2})$ is positive on $\mathbf{X}_{\operatorname{Re}}$;
  3. 3.
    the quadratic module $\mathcal{T}^{\mathbb{R}}(\mathbf{X}_{\operatorname{Re}})$ is Archimedean.
The first assumption is true because the original problem has a non-empty feasible set $\mathcal{Q}(\rho,\nu)$: for any spectral decompositions $\rho=\sum_{\ell}\omega_{\ell}u_{\ell}u_{\ell}^{*}$ and $\nu=\sum_{j}\omega^{\prime}_{j}v_{j}v_{j}^{*}$, one has $\{(\omega_{\ell}\omega^{\prime}_{j},u_{\ell},v_{j})_{\ell,j}\}\in\mathcal{Q}(\rho,\nu)$.
The second assumption holds with $\Lambda_{1}=\Gamma_{1}=I_{n}$ and $\Lambda_{2}=\Gamma_{2}=0$, since the polynomial $\operatorname{Tr}((\mathbf{a}\mathbf{a}^{T}+\mathbf{b}\mathbf{b}^{T})+(\mathbf{c}\mathbf{c}^{T}+\mathbf{d}\mathbf{d}^{T})=\|\mathbf{a}\|+\|\mathbf{b}\|+\|\mathbf{c}\|+\|\mathbf{d}\|$ is equal to $2>0$ on $\mathbf{X}_{\operatorname{Re}}$.
The third assumption is obviously true since the polynomial $2-\|\mathbf{a}\|-\|\mathbf{b}\|-\|\mathbf{c}\|-\|\mathbf{d}\|$ belongs to $\mathcal{T}^{\mathbb{R}}(\mathbf{X}_{\operatorname{Re}})$.
The result from Lemma 10 can then be applied together with Theorem 3 to obtain that for all $t\geq 4nd\sqrt{d}=32n$, one has
$0\leq W^{2}_{2}(\rho,\nu)-W^{2}_{2}(\rho,\nu)_{2t}\leq\frac{\kappa(n)}{t^{2}},$
$\text{with }\kappa(n):=\frac{3}{2}C_{\mathbf{X}_{\operatorname{Re}}(2n,4)}\left(3f_{\operatorname{Re},\max}+\|w^{\operatorname{opt}}\|_{1}h_{\max}\right),$ (49)
where
  • the numerator of the constant $\frac{3}{2}$ has been obtained by adding 1 to the evaluation of the dual objective function at $\Gamma_{1}=\Lambda_{1}=I_{n}$ and $\Lambda_{2}=\Gamma_{2}=0$, that is the tuple we used above to satisfy the last assumption from Lemma 10. The denominator corresponds to the minimum of $\operatorname{Tr}((\mathbf{a}\mathbf{a}^{T}+\mathbf{b}\mathbf{b}^{T})+(\mathbf{c}\mathbf{c}^{T}+\mathbf{d}\mathbf{d}^{T})=\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}+\|\mathbf{c}\|^{2}+\|\mathbf{d}\|^{2}$ on $\mathbf{X}_{\operatorname{Re}}$;
  • the constant $f_{\operatorname{Re},\max}$ is the maximum of $f_{\operatorname{Re}}$ on $\mathbf{X}_{\operatorname{Re}}$;
  • the constant $\|w^{\operatorname{opt}}\|_{1}$ is the sum of absolute values of all entries of a tuple $(\Lambda_{1},\Lambda_{2},\Gamma_{1},\Gamma_{2})$ achieving the supremum in the dual (48);
  • the constant $h_{\max}$ is the maximum value over $\mathbf{X}_{\operatorname{Re}}$ taken by the entries of the four matrix polynomials involved in the equality constraints of the primal (47).
We end the proof by providing estimates for the 3 quantities $f_{\operatorname{Re},\max}$, $\|w^{\operatorname{opt}}\|_{1}$ and $h_{\max}$.
1. The first quantity $f_{\operatorname{Re},\max}$ is equal to the maximum $f_{\max}$ of $f$ on $\mathbf{X}$. One has
$f(\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}})=\operatorname{Tr}[(\mathbf{x}\mathbf{x}^{*}-\mathbf{y}\mathbf{y}^{*})\overline{(\mathbf{x}\mathbf{x}^{*}-\mathbf{y}\mathbf{y}^{*})}]$ $=\left|\sum_{i=1}^{n}x_{i}^{2}\right|^{2}+\left|\sum_{i=1}^{n}y_{i}^{2}\right|^{2}-2\left|\sum_{i=1}^{n}x_{i}y_{i}\right|^{2}$
$\leq\left(\sum_{i=1}^{n}|x_{i}|^{2}\right)^{2}+\left(\sum_{i=1}^{n}|y_{i}|^{2}\right)^{2}.$
This shows that $f_{\max}\leq 2$. In addition, $f(\mathbf{x},\overline{\mathbf{x}},\mathbf{y},\overline{\mathbf{y}})=2$ for $\mathbf{x}=(1,0,\dots,0)$ and $\mathbf{y}=(0,1,0,\dots,0)$. Therefore, one has $f_{\max}=f_{\operatorname{Re},\max}=2$.
2. Since the complex dual (38) is attained at a bounded tuple $(\Lambda,\Gamma)$, the real dual (48) is also attained at a bounded tuple $(\Lambda_{1},\Lambda_{2},\Gamma_{1},\Gamma_{2})$, with $\Lambda=\Lambda_{1}+\mathbf{i}\,\Lambda_{2}$ and $\Gamma=\Gamma_{1}+\mathbf{i}\,\Gamma_{2}$.
For all $z=a+\mathbf{i}\,b$ with $a,b\in\mathbb{R}$, one has $|a|+|b|\leq\sqrt{2}\sqrt{a^{2}+b^{2}}=\sqrt{2}|z|$, thus $\|\Lambda_{1}\|_{1}+\|\Lambda_{2}\|_{1}\leq\sqrt{2}\|\Lambda\|_{1}$, and $\|\Gamma_{1}\|_{1}+\|\Gamma_{2}\|_{1}\leq\sqrt{2}\|\Gamma\|_{1}$.
3. The third quantity $h_{\max}$ is equal to 2, since each term of the form $a_{i}a_{j}+b_{i}b_{j}$, $a_{i}b_{j}-b_{j}a_{i}$, $c_{i}c_{j}+d_{i}d_{j}$, or $c_{i}d_{j}-d_{j}c_{i}$, is upper bounded by 2 on $\mathbf{X}_{\operatorname{Re}}$.
We have just proved that
$\kappa(n)\leq\frac{3}{2}\left(3+\sqrt{2}(\|\Lambda\|_{1}+\|\Gamma\|_{1})\right)C_{\mathbf{X}_{\operatorname{Re}}(2n,4)}.$ (50)
5 Concluding remarks
We have proved a convergence rate in $O(1/t^{2})$ for the Schmüdgen-type (or equivalently Putinar-type) hierarchies of lower bounds $\operatorname{lb}(q,\mathcal{T}(\mathbf{X}))_{2t}$ for the minimization of a polynomial over the bi-sphere. Thanks to this result, we could provide a similar rate for a hierarchy of semidefinite programs approximating the $2$ order quantum Wasserstein distance. We also extended this result to arbitrary sphere products, and products of distinct sets while assuming that a rate is available for each set. Our proof technique heavily relies on the polynomial kernel method.
We would like to emphasize that all obtained convergence rates for hierarchies of lower bounds also hold for the hierarchies of upper bounds developed in [Las11], by following the same reasoning as the one from [Slo22, Section 5]. In addition, our derived rates can be combined with the general result from [GM25] to obtain rates for GMP instances with measures supported on set products.
Recently, the convergence rate of the Putinar-type hierarchy for polynomial optimization over the hypercube has been revisited in [GdKV25]. The idea is also to rely on the polynomial kernel method with a suitably chosen Gaussian density. An interesting open question would be to investigate whether this idea could be extended for set products.
For the case of the unit sphere, we would like also to mention the recent works [LJ23, BL24], where the authors provide convergence rates in $O(1/t)$ for a specialized hierarchy of spectral bounds, using quantum and real de Finetti theorems, respectively. A natural research track would be to extend this analysis to the case of the bi-sphere.
Eventually, it has been proved in [Nie14] that the Putinar-type hierarchy converges in finitely many steps, under genericity assumptions on the objective function and the polynomials describing the feasible set. When optimizing a generic polynomial over the unit sphere, it has been shown in [Hua23] that one also has finite convergence. Here again, it would be interesting to obtain similar finite convergence results for specific set products such as the bi-sphere.
Acknowledgments.
We thank Jonas Britz, Saroj Prasad Chhatoi, Monique Laurent and Lucas Slot for fruitful discussions. This work was supported by the European Union’s HORIZON–MSCA-2023-DN-JD programme under the Horizon Europe (HORIZON) Marie Sklodowska-Curie Actions, grant agreement 101120296 (TENORS).